What is the probability of not drawing a green marble from a jar containing 7 red, 4 green, 3 white, and 9 blue marbles?
\(7+4+3+9\)
\(11+12\)
\(23\)
\(\frac{23}{23} - \frac{4}{23}\)
\(\frac{23-4}{23}\)
\(\frac{19}{23}\) \(chance\)
Answer:
Probability of not drawing a green marble is 19/23.
Step-by-step explanation:
A jar is containing 7 red, 3 white, and 9 blue marbles.
Then the probability to draw a green marble will be = Numbers of green marbles/ Total numbers of marbles of all colors
P(G) = 4/(7 + 4 + 3 + 9)
P(G) = 4/23
Now the probability of not drawing a green marble = 1 - probability of drawing a green ball
P = 1 - 4/23 = 19/23
Therefore answer is 19/23.